OpenStudy (anonymous):

can some one help me with this solving this equation that has rational expressions x-3/x-1-2x-4/x-1=0

OpenStudy (anonymous):

i do not think that there is a solution i keep coming up with 0 i just want to make sure that is correct

OpenStudy (anonymous):

if you multiply everything under x-3 you are left with x-3 = 0 so x could be 3... as long as x isn't -5 or 1 because you don't want to divide be zero. so one possible answer is 3. That's the best I can do right now...There seem to be various ways of manipulating but then you run into division by zero..

OpenStudy (anonymous):

I say the answer is 3 but don't quote me on that..

OpenStudy (anonymous):

x-3/x-1-2x-4/x-1=0 is this like this (x-3)/(x-1)-(2x-4)/(x-1)=0 ? therefore [(x-3)-(2x-4)]/(x-1)=0 -x-3+4=0(x-1) -x+3=0 x=3

OpenStudy (gw2011):

If the problem is: (x-3)/[(x-1-2x-4)/(x-1)]=0, then it becomes: [(x-3)(x-1)]/(x-1-2x-4)=0 [(x-3)(x-1)]/(-x-5)=0 (x-3)(x-1)=0 x-3=0 x=3 x-1=0 x=1 If you substitute x=3 and x=1 into [(x-3)(x-1)]/(-x-5), then there is no division by zero and the answer equals zero for both values of x.